OpenStudy (anonymous):

M/M-5 - 3/M+5= 2/M^2-25

OpenStudy (anonymous):

\[m^2-25\]is the same as \[(m+5)(m-5)\], so multiply both sides by that and you get \[m(m+5)-3(m-5)=2\]Just factor everything out and solve the quadratic equation.

OpenStudy (dumbcow):

no real solution use quadratic formula, but you will get neg num inside sqrt sqrt(-1) = i

OpenStudy (anonymous):

-33

OpenStudy (anonymous):

MR.COW AND ARXANAS I CAME UP WITH -33 WHAT DO YOU TOO THINK

OpenStudy (dumbcow):

m=-33 ? no thats not right m= [-2 +-sqrt(4-(4*13))]/2

OpenStudy (anonymous):

SQUARE ROOT35

OpenStudy (dumbcow):

sqrt(4-52) = sqrt(-48) = sqrt(48)*i

OpenStudy (anonymous):

ITS 3

OpenStudy (dumbcow):

answer: m = -1+-2sqrt(3)*i